A horizontal force F applied to an object varies with its displacement. (Force is 4.0) Calculate the work done by the force when the object moves from x =0 to x=10.0m.

1
Expert's answer
2021-12-06T09:49:35-0500

By the definition of the work done, we have:


"W=Fd."

Let's first find the displacement of the object:


"d=x_f-x_i=10\ m-0\ m=10\ m."

Finally, we can find the work done by the force:


"W=4\ N\times 10\ m=40\ J."